### v2

# Federated EndoViT: Pretraining Vision Transformers via Federated Learning on Endoscopic Image Collections

Abstract
Purpose: In this study, we investigate the training of foundation models using federated learning to address data-sharing limitations and enable collaborative model training without data transfer for minimally invasive surgery. Methods: Inspired by the EndoViT study, we adapt the Masked Autoencoder for federated learning, enhancing it with adaptive Sharpness-Aware Minimization (FedSAM) and Stochastic Weight Averaging (SWA). Our model is pretrained on the Endo700k dataset collection and later fine-tuned and evaluated for tasks such as Semantic Segmentation, Action Triplet Recognition, and Surgical Phase Recognition. Results: Our findings demonstrate that integrating adaptive FedSAM into the federated MAE approach improves pretraining, leading to a reduction in reconstruction loss per patch. The application of FL-EndoViT in surgical downstream tasks results in performance comparable to CEN-EndoViT. Furthermore, FL-EndoViT exhibits advantages over CEN-EndoViT in surgical scene segmentation when data is limited and in action triplet recognition when large datasets are used. Conclusion: These findings highlight the potential of federated learning for privacy-preserving training of surgical foundation models, offering a robust and generalizable solution for surgical data science. Effective collaboration requires adapting federated learning methods, such as the integration of FedSAM, which can accommodate the inherent data heterogeneity across institutions. In future, exploring FL in video-based models may enhance these capabilities by incorporating spatiotemporal dynamics crucial for real-world surgical environments.
1 Graphical Abstract
2 Introduction
The goal of Surgical Data Science (SDS) is to assist surgeons by recording, analyzing, and modeling surgical data [1]. AI introduced analysis of laparoscopic video streams has the potential to aid navigation and decision-making, improve surgical outcomes, reduce morbidity, and enable surgeons to learn from recorded procedures [2]. To develop such AI-driven assistance systems, robust and powerful models are required to accurately interpret surgical data and provide meaningful insights [2]. It has been shown that Foundation Models are demonstrably robust and powerful models [3]. Bommasani et al. [3] define Foundation Models as follows: “A foundation model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks […].” A foundation model for SDS is, for example, the Endoscopic Vision Transformer (EndoViT) [4]. This model is pretrained on 700,000 surgical images and has demonstrated strong performance on surgical tasks like Surgical Phase Recognition (SPR), Action Triplet Recognition (ATR), and Surgical Scene Segmentation (SSS). Moreover, the pretrained EndoViT model is already providing meaningful representation when the model is fine-tuned on a limited amount of data.
However, a foundation model trained on limited or homogeneous data may fail to generalize across different patient groups, hospitals, or procedures [5]. In this context, homogeneous data refers to datasets collected from a single institution, patient population, or surgical procedure, lacking the variability needed for robust generalization. This variability stems from differences in patient anatomy, surgical techniques, equipment, and clinical protocols [5], making the development of foundation models for surgical applications particularly challenging. Moreover, data protection regulations such as the Health Insurance Portability and Accountability Act (HIPAA) and the General Data Protection Regulation (GDPR) impose strict constraints on patient data sharing, making cross-institutional data aggregation legally and organizationally complex. Federated Learning (FL) offers a solution for this predicament by enabling collaborative model training across medical institutions without data sharing, maintaining privacy while leveraging diverse datasets [6, 3]. FL faces challenges from heterogeneous data, e.g., label, feature, quantity, and quality skew [7], which must be addressed for effective model convergence and generalization. Otherwise federated training fails in comparison to centralized training.
Despite the early success of FL in surgical tasks, as evidenced by Kassem et al. [8] for SPR, there is a significant gap in the development of foundation models tailored for surgical applications. In contrast, recent work on adaptive federated knowledge injection into medical MRI and CT foundation models [9], as well as FL in real-world medical imaging scenarios [10], highlights the potential for scaling such models in other medical domains. However, both works do not yet address surgical data specifically. AI models for SDS trained via FL remain largely task-specific. One reason for this may be the limited availability of large-scale, multicentric surgical datasets. Another important factor is that surgical data tends to be highly heterogeneous and less standardized compared to radiological imaging data. A compelling example of this heterogeneity is the Critical View of Safety (CVS) challenge, which aggregates cholecystectomy videos from 67 surgeons across 57 countries [11], illustrating the substantial variability in surgical technique, recording conditions, and clinical environments.
Those limitations raise the question of how federated learning can be leveraged to train a foundation model that generalizes across diverse, heterogeneous surgical datasets and tasks while preserving privacy, and ensuring that federated learning performs on par with centralized training.
Contributions: By leveraging federated learning, this work introduces a foundation model specifically designed for the surgical domain, enabling collaborative model development without the need for data sharing. To this end, we extend the EndoViT framework for federated training and integrate adaptive Federated Sharpness-Aware Minimization as proposed by Caldarola et al. [12]. Our approach addresses critical challenges in SDS related to data privacy, institutional heterogeneity, and limited data accessibility, achieving performance comparable to centralized training. This demonstrates that federated foundation models can offer a privacy-preserving and robust solution for SDS, with effective collaboration requiring the adaptation of suitable federated learning methods, such as the integration of FedSAM, which accommodates the inherent inter-center data heterogeneity.
3 Related Work
This work builds on the results of EndoViT by Batić et al. [4] and Sharpness-Aware Minimization (SAM) of Caldarola et al. [12]. Therefore, we first introduce the key concepts of the EndoViT study and SAM before detailing our adaptation to a FL setting.
3.1 EndoViT as Centralized Baseline
The EndoViT model is composed of two stages. Initially, EndoViT is pretrained in a self-supervised learning (SSL) fashion. Subsequently, the trained backbone of the EndoViT model is utilized as an encoder and fine-tuned for a variety of downstream tasks in the second stage.
As first step the EndoViT model, a Vision Transformer (ViT), was pretrained on the Endo700k dataset collection using the Masked Autoencoder (MAE) approach proposed by He et al. [13]. The MAE method is a SSL technique that reconstructs the missing components of the input images. Specifically, the image is initially divided into 256 non-overlapping patches, where 75 % of these patches are masked randomly. Subsequently, the encoder is trained to process the visible patches to generate a visual representation. A lightweight decoder is then employed to reconstruct the masked patches. This MAE approach enables the effective learning of meaningful representations. Mean-squared error (MSE) is used to calculate the reconstruction loss.
In the second phase, the performance of the encoder is tested on three downstream tasks.
For Surgical Scene Segmentation (SSS), the goal is to assign each pixel in an image to a predefined class. A Dense Prediction Transformer (DPT) [14] is proposed, which uses a Vision Transformer (ViT) backbone instead of traditional convolutional layers, paired with a convolutional decoder. This architecture, which avoids explicit down-sampling, preserves higher resolution and enables accurate, globally coherent predictions. CholecSeg8k [15] is used as dataset for this downstream task. As metric, the Intersection over Union (IoU) is reported for all SSS experiments.
Action Triplet Recognition (ATR) involves identifying the relevant tool, action, and target organ from an image presented as a triplet <instrument, verb, target >. For this task, the authors recommended a model in EndoViT with a feature extraction backbone and a linear head. The evaluation is conducted exclusively on the fifth fold of the ATR split of the CholecT45 dataset [16] to ensure a fair and representative performance assessment. This particular fold is selected due to its optimal balance in class distribution [4, 16]. For ATR the mean Average Precision (mAP) is reported.
In Surgical Phase Recognition (SPR), the goal is to classify each video frame into predefined surgical phases. In the EndoViT study, the TeCNO model [17] is used, which combines Multi-Stage Temporal Convolutional Networks (MS-TCNs) with the EndoViT encoder as a backbone. The large temporal receptive field of TeCNO is particularly advantageous for efficient training and inference on long, unedited surgical videos. Cholec80 [18] is the dataset used for this downstream task. The metric used for SPR is accuracy (acc).
In addition to fine-tuning on all available data, the EndoViT study introduces fine-tuning with varying amounts of labeled data. These experiments are designated as few-shot tuning.
The results show that the EndoViT model matches or exceeds the performance of the ImageNet pre-trained ViT model in the previously specified downstream tasks. In particular, surgical domain-specific pretraining yields significant improvements in complex tasks such as ATR or SSS, where the ImageNet-pretrained models are outperformed. For simpler tasks, such as SSS, the benefits of domain-specific pretraining are less pronounced, but still do not worsen the results. These results suggest that EndoViT serves as an effective feature extraction backbone for surgical image understanding, especially when dealing with limited amounts of annotated data.
3.2 Adaptive Federated Sharpness-Aware Minimization
Caldarola et al. [12] address the challenges posed by degraded performance and limited generalization in FL settings where data is heterogeneous. The authors argue that this problem is linked to the sharpness of the loss landscape, where sharper minima correlate with poorer generalization. They propose two strategies to mitigate this.
One strategy introduces Sharpness-Aware Minimization (SAM) on the client side. SAM modifies training to target flatter regions in the loss landscape by minimizing both the loss value and its sharpness. Flat regions are areas where small perturbations in the weights lead to only minor increases or decreases in the loss. By seeking parameters that lie within neighborhoods of uniformly low loss, SAM guides the model toward flatter minima. This approach makes the model less sensitive to data variations, thereby enhancing its generalization across diverse client distributions.
Secondly, they employ Stochastic Weight Averaging (SWA) on the server to aggregate client models. SWA involves averaging weights over multiple training epochs, promoting convergence to flatter minima. As demonstrated by Izmailov et al. [19], this process tends to identify broader minima in the loss landscape, which are associated with superior generalization. This is due to the averaging of weights over time, which helps to mitigate sharp minima and consequently leads to a global model that exhibits superior generalization to unseen data.
Empirical evaluations demonstrate that the integration of both approaches results in substantial enhancements in model performance across a range of vision tasks, including large-scale classification, semantic segmentation, and domain generalization.
4 Methodology
4.1 Federated EndoViT
In this subsection, the contributions and adjustments made to the EndoViT model for FL are described. The study is comprised of two primary experiments, as illustrated in Figure 1: pretraining and fine-tuning on downstream tasks. Firstly, a federated variant of the EndoViT model, designated as FL-EndoViT, is pretrained. Secondly, we assess the efficacy of the trained FL-EndoViT backbone across SSS, ATR, and SPR, and compare it to the centralized results of EndoViT, named CEN-EndoViT.
Figure 1: Federated EndoViT Framework: A two-stage approach for training a surgical Foundation Model using Federated Learning. (I) The model undergoes federated self-supervised pretraining with a Masked Autoencoder strategy, where 75% of the 256 image patches are masked and reconstructed. Adaptive Federated Sharpness-Aware Minimization is applied on the client side, while Stochastic Weight Averaging is used on the server side to enhance generalization across centers. (II) The pretrained EndoViT encoder serves as a Foundation Model, enabling fine-tuning and feature extraction for diverse surgical downstream tasks. This approach builds upon the work of Batić et al. [4].
Self-Supervised Pretraining
As shown in Figure 1 (I) the initial phase, designated as self-supervised pretraining, is executed within the framework of a FL environment. A standard FL configuration consists of multiple decentralized centers that collaboratively train a shared global model, with model updates being aggregated by a common server. Each center retains its own local dataset, which is kept private.
Each FL-EndoViT center is performing a local training round where the objective of the pretraining is to reconstruct the original input image. For this the input is split into 256 non-overlapping patches, with 75% of the patches being masked out and fed to the EndoViT model. In comparison to CEN-EndoViT, we employ the adaptive FedSAM algorithm as the local optimizer on each client. This approach is designed to address the challenge of sharp minima, which arises due to heterogeneous, non-identically distributed (non-I.I.D.) client data. It is anticipated that this methodology will enhance generalization, convergence, and reduce the risk of over-fitting. Following a local training round, the EndoViT models are transmitted to a central server, where they are aggregated using the federated averaging (FedAvg) algorithm.
Furthermore, we employ SWA, which is also utilized in CEN-EndoViT. This process involves the averaging of weights from disparate training epochs, thereby ensuring that solutions are confined to the regions of the loss landscape that are both smooth and stable. The implementation of this process on the server side during the final five epochs is particularly advantageous, as it enables the aggregation of diverse client updates from multiple rounds.
Fine-tuning on Downstream Tasks
In the second phase, as shown in Figure 1 (II) fine-tuning on downstream tasks is conducted in a centralized manner, with the pretrained FL-EndoViT model used as an encoder across the three tasks.
A prevailing trend in Foundation Models is to maintain the frozen state of the model [20, 21, 22]. This practice is driven by numerous factors. One such reason is the potential to reduce training costs and accelerate convergence. Additionally, a pretrained backbone captures general surgical features. Freezing the backbone ensures that task-specific training does not overwrite essential features and avoids overfitting, thereby maintaining the backbone’s robustness and generalizability. To this end, we have conducted a series of ablation experiments in which the backbone is maintained in a frozen state. These experiments are designed to ascertain the potential for enhanced efficiency in the utilization of both FL-EndoViT and CEN-EndoViT.
In addition to reporting acc for SPR, we also provide the F1-score. As discussed by Funke et al. [23], relying solely on accuracy can be misleading when evaluating classification performance on imbalanced datasets like Cholec80, as it doesn’t account for unequal class distributions and may lead to overly optimistic assessments.
4.2 Implementation Details
The authors of EndoViT have combined nine publicly available surgical datasets into the Endo700k dataset collection [4] with approximately 700,000 images. Each dataset consists of raw minimally invasive surgery (MIS) images from a range of different procedures such as prostatectomy, cholecystectomy, gastrectomy, proctocolectomy, rectal resection and sigmoid resection performed both laparoscopically and robotically.
To prevent overlap between pretraining and fine-tuning datasets, three variants of Endo700k were created, as Cholec80 was used for pretraining while fine-tuning involved CholecT45, Cholec80, and CholecSeg8k. Therefore, any images that appear in their validation or test sets are excluded from the pretraining dataset resulting in Endo700k for SSS, Endo700k for ATR, and Endo700k for SPR.
In our FL setting, each client has access to one of the nine sub-datasets of Endo700k, while the server operates without direct access to any data.
We implemented our training procedures using the Flower FL Framework [24] in simulation mode, with PyTorch 1.13.0. In our federated setup, each of the nine Endo700k datasets serves as a distinct center. Each center stores its best local model, and the server maintains the best averaged global model. Pretraining in simulation mode was carried out in parallel on four NVIDIA V100 GPUs that do not share memory, while the centralized baseline was retrained on a single NVIDIA V100 GPU. Centralized fine-tuning for SSS, ATR, and SPR was performed on a NVIDIA V100 GPU. Hyperparameters were adopted from the EndoViT Git [4] repository to ensure comparability. Our code is available under following link.
5 Experiments
This study involved a series of experiments, beginning with an ablation study to show that integrating FedSAM into the MAE pretraining is critical for achieving performance on par with centralized training. Following this, we conducted a comparative analysis of FL-EndoViT and CEN-EndoViT across three clinically relevant surgical downstream tasks. This process requires a fine-tuning procedure to optimize the performance of these models on the specific tasks. This investigation aims to ascertain the need of fine-tuning the backbone entirely, or whether fine-tuning the task-specific decoder, with the backbone frozen, is sufficient. To investigate this, we carried out a second ablation study. Finally, a comparison experiment on full-shot and few-shot training was conducted, following the clarification of the EndoViT model fine-tuning procedure. To ascertain the statistical significance of the discrepancy in performance between the two models, we employed a paired Wilcoxon signed-rank test. For SSS, the test is performed at the image level, whereas for the other two downstream tasks, it is conducted at the video level. The null hypothesis posits that the two models perform similarly. The significance level, denoted by alpha, is set at 0.01. The significance level, denoted by alpha, is set at 0.01. This is done so that stronger evidence is required before the null hypothesis can be rejected, thereby emphasizing the significance of the results obtained.
5.1 Federated Pretraining Ablation Study
The first stage of EndoViT is pretraining. We conducted an ablation study to evaluate whether integrating adaptive FedSAM enhances performance compared to a standard FL approach. To do so, we trained three models per downstream task: CEN-EndoViT, FL-EndoViT without adaptive FedSAM, and FL-EndoViT with adaptive FedSAM. Pretraining was performed separately for each downstream task using the task-specific version of the Endo700k dataset. As outlined in the methods, all three EndoViT variants incorporate SWA, resulting in an additional SWA-enhanced model for each variant.
To assess models reconstruction performance, we analyzed the number of image patches with a Mean Squared Error (MSE) reconstruction loss below predefined thresholds: 0.3, 0.1, 0.05, and 0.01. During self-supervised pretraining, EndoViT reconstructs masked image patches, and the quality of this reconstruction reflects the model’s ability to learn meaningful visual representations. A higher number of patches below a given threshold indicates more precise reconstruction, preserving critical structures and details essential for downstream surgical tasks. Notably, lower thresholds correspond to higher confidence in reconstruction, while higher thresholds may still include reconstruction noise.
Our results, shown in Table 1, indicate that the federated MAE approach alone underperforms across all three EndoViT models, each pretrained for a different downstream task in the federated setting. Integrating adaptive FedSAM reduces reconstruction error and increases the number of accurately reconstructed patches across different thresholds. However, while it improves upon standard federated training, it does not fully match the performance of the centralized model, though it remains competitive.
While the SWA-enhanced model does not surpass the standard model in validation metrics, it remains valuable due to its potential for improved generalization. Caldarola et al. [12] suggests that SWA smoothens the optimization landscape by averaging model weights over multiple epochs, allowing convergence toward flatter minima and improving robustness to data variations. This effect is particularly relevant in FL, where heterogeneous data poses a challenge. Additionally, SWA improves stability and consistency across datasets. Previous studies indicate that combining adaptive FedSAM with SWA enhances data augmentation, leading to flatter minima and potentially better generalization than adaptive FedSAM alone. However, our results do not provide conclusive evidence of superior generalization in this specific setup.
Based on these findings and insights from the literature, we select the FL-EndoViT model with adaptive FedSAM and SWA for further evaluation. In the next section, we compare its performance against CEN-EndoViT with SWA to determine whether the FL backbone is as effective as its centralized counterpart.
Table 1: Pretraining Performance Comparison: The distribution of patch count below different thresholds of the reconstruction loss for Surgical Scene Segmentation (SSS), Action Triplet Recognition (ATR), and Surgical Phase Recognition (SPR), shows the importance of combining MAE with adaptive FedSAM to get results that correspond to centralized (CEN) pretraining.
Maximum Number of Patches below Threshold
Pretraining for SSS ATR SPR
Variant CEN MAE+SAM MAE CEN MAE+SAM MAE CEN MAE+SAM MAE
Patches below 0.3 171 168 67 172 168 67 172 167 67
Patches below 0.3 SWA 171 166 67 172 167 67 172 165 67
Patches below 0.1 126 112 25 127 112 25 127 110 25
Patches below 0.1 SWA 120 108 25 121 110 25 124 105 25
Patches below 0.05 92 82 11 96 81 11 96 80 11
Patches below 0.05 SWA 90 80 11 91 80 11 92 77 11
Patches below 0.01 45 39 0 45 39 0 46 36 0
Patches below 0.01 SWA 45 39 0 46 39 0 46 36 0
5.2 Downstream Tasks
The goal of SSL pretraining is to develop encoders that serve as backbones for various downstream tasks. Consequently, we evaluate the superior variant of the FL-EndoViT backbone with adaptive FedSAM and SWA from the previous section on three clinically relevant tasks: SSS, ATR, and SPR.
Our downstream task evaluation is divided into two main parts. First, we conduct an ablation study to determine whether full fine-tuning is necessary or if freezing the backbone and fine-tuning only the task-specific head is sufficient. Second, we analyze the better-performing variant in detail, extending the evaluation to include fine-tuning on all available data as well as few-shot learning, where models are trained on a limited number of videos.
All experiments are conducted in a centralized environment. To ensure robustness and reliability, we perform three independent technical replications. For few-shot experiments, we use different sets of videos across runs to further assess model consistency. The results reported in this paper represent the average over all technical replications.
This experimental setup allows for a direct comparison between the performance of the FL-EndoViT and its centralized counterpart CEN-EndoViT.
5.2.1 Ablation Study: Fine-tuning vs. Frozen Backbone
Table 2: Impact of Frozen Fine-Tuning in comparison to Fully Fine-Tuning: This table compares the performance of fully fine-tuned and frozen-backbone models across Surgical Scene Segmentation (SSS), Action Triplet Recognition (ATR), and Surgical Phase Recognition (SPR). The reported values represent the mean metric computed per image or video, followed by the mean and standard deviation calculated across all images. The bold values indicate whether CEN-EndoViT or FL-EndoViT performs better. Statistically significant differences, determined using the Wilcoxon signed-rank test with $\alpha=0.01$, are marked with an asterisk (*).
Task SSS ATR
High Res Low Res
Metric IOU mAP
Full CEN-EndoViT $\bm{67.81\%\pm 8.03\%}$* $\bm{67.54\%\pm 8.06\%}$* $31.33\%\pm 4.61\%$
FL-EndoViT $65.98\%\pm 8.10\%$ $67.42\%\pm 8.18\%$ $\bm{40.79\%\pm 6.65\%}$*
Frozen CEN-EndoViT $\bm{68.57\%\pm 8.14\%}$* $\bm{67.28\%\pm 8.94\%}$* $\bm{31.33\%\pm 4.61\%}$*
FL-EndoViT $67.23\%\pm 7.37\%$ $65.22\%\pm 7.55\%$ $26.90\%\pm 5.58\%$
Task SPR
S1 S2 S1 S2
Metric ACC F1
[…]
Frozen CEN-EndoViT $\bm{72.26\%\pm 8.96\%}$* $\bm{79.88\%\pm 10.88\%}$* $\bm{59.71\%\pm 8.70\%}$* $\bm{75.12\%\pm 9.50\%}$*
FL-EndoViT $65.72\%\pm 10.72\%$ $75.26\%\pm 11.55\%$ $53.39\%\pm 8.82\%$ $72.29\%\pm 9.46\%$
An ablation study was conducted on three tasks—SSS, ATR, and SPR—to evaluate whether full fine-tuning of the EndoViT model is required, or if updating only the task-specific head while keeping the backbone frozen is sufficient. We trained models on the full dataset and compared the performance of fully fine-tuned models against those with a frozen EndoViT backbone. The performance differences are summarized in Table 2.
Our results indicate that freezing the EndoViT backbone and fine-tuning only the task-specific head generally leads to suboptimal performance across most tasks. The degradation is particularly severe when FL-EndoViT is used, especially for ATR and SPR, where full fine-tuning yields substantial improvements, up to 13.89% for ATR, 17.80 % for SPR-S1 (F1), and 12.76% for SPR-S2 (F1). For CEN-EndoViT, the performance gap is less pronounced but still noticeable. Interestingly, in specific cases such as SSS-High Res, freezing the backbone even results in slight performance improvement, suggesting that important pretrained representations are lost when the backbone is not remained fixed. Nevertheless, the observed gap in centralized and federated settings, full fine-tuning is beneficial for extracting task-specific representations.
Given these findings, we focus exclusively on the fully fine-tuned experiments in the subsequent sections. Full fine-tuning outperforms the frozen-backbone approach across nearly all tasks, particularly for FL-EndoViT where freezing the backbone results in significant performance loss. By concentrating on fully fine-tuned models, we ensure that our analysis reflects the most effective approach for leveraging the EndoViT backbone in SDS applications.
5.2.2 Fine-tuning Experiment
We evaluate on all three downstream tasks from the EndoViT paper by fully fine-tuning the model, which includes the EndoViT backbone and a task-specific head, in the first fine-tuning experiment.
Table 3: Impact of Few-Shot Fine-Tuning on Surgical Scene Segmentation (SSS): This table compares the SSS performance of FL-EndoViT and CEN-EndoViT in a few-shot setting across different training set sizes. The reported values represent the mean IoU-Score computed per image over the technical repetitions, followed by the mean and standard deviation caclucation across all images. The bold values indicate weather CEN-EndoViT or FL-EndoViT performs better. Statistically significant differences, determined using the Wilcoxon paired test with $\alpha=0.01$, are marked with an asterisk (*).
Fine-Tuned on Res. CEN-EndoViT FL-EndoViT
1 Video Low $\bm{42.87\%\pm 6.84\%}$* $40.50\%\pm 7.60\%$
High $25.68\%\pm 8.32\%$ $\bm{42.03\%\pm 6.50\%}$*
2 Videos Low $\bm{56.60\%\pm 5.54\%}$* $53.32\%\pm 5.40\%$
High $37.12\%\pm 3.82\%$ $\bm{56.39\%\pm 6.39\%}$*
4 Videos Low $\bm{61.06\%\pm 5.62\%}$* $58.98\%\pm 6.25\%$
High $45.51\%\pm 5.61\%$ $\bm{60.34\%\pm 6.58\%}$*
Surgical Scene Segmentation
The results of the SSS fine-tuning experiment are detailed in Table 3, while the outcomes of comprehensive fine-tuning on the entire dataset are reported in Table 2.
In the low-resolution setting, the results demonstrate that the model with the CEN-EndoViT backbone generally outperforms the FL-EndoViT model. The centralized model achieved significantly higher IoU scores than the FL model when trained on one, two, or four videos. However, as the training data increases towards all videos used for fine-tuning, the performance gap narrows, with fine-tuned FL model achieving comparable results.
This trend is reversed in the high-resolution setting, where the model with the FL-EndoViT backbone consistently outperforms the centralized model in all settings. The advantage of using the FL backbone is particularly pronounced in smaller training sets, where it achieves significantly higher IoU scores. For instance, the model with FL-EndoViT achieved an average mIoU of 42.03% for a single video, compared to centralized’s 25.68%, marking a significant improvement. A similar enhancement is observed in the performance of centralized models when trained on two videos, which exhibits an increase of 19.27%. Meanwhile, models with FL-EndoViT and CEN-EndoViT exhibit only a small performance gap for larger training sets, such as when trained on all available videos, where CEN-EndoViT outperforms the federated setting by 1.83%. These results highlight the effectiveness of FL in scenarios where data availability is limited.
Action Triplet Recognition
Table 4: Impact of Few-Shot Fine-Tuning on Action Triplet Recognition (ATR): This table compares the ATR performance of FL-EndoViT and CEN-EndoViT in a few-shot setting across different training set sizes. The reported values represent the mean average precision score computed per image over the technical repetitions, followed by the mean and standard deviation caclucation across all images. The bold values indicate weather CEN-EndoViT or FL-EndoViT performs better. Statistically significant differences, determined using the Wilcoxon paired test with $\alpha=0.01$, are marked with an asterisk (*).
Fine-Tuned on CEN-EndoViT FL-EndoViT
[…]
4 Videos $26.00\%\pm 4.18\%$ $\bm{27.24\%\pm 4.75\%}$
[…]
The results of the ATR fine-tuning experiment are presented in Table 4, while the outcomes of full fine-tuning on the entire dataset are provided in Table 2.
The model employing the FL backbone demonstrates superior performance in most cases when evaluated against the mean average. In scenarios where the FL model is trained on the complete training dataset, it attains the highest mAP across all iterations, exhibiting a substantial enhancement over the centralized variant. For instance, the model utilizing the FL backbone attains a mAP of 40.79%, in comparison to the centralized model’s 31.33%, signifying a notable improvement. However, in the few-shot setting, the performance gap between models with the centralized and FL backbone becomes negligible, suggesting that the federated approach can achieve comparable results with significantly less data. This highlights the strong generalization ability of the FL-trained backbone in low-data regimes.
Surgical Phase Recognition
Table 5: Impact of Few-Shot Fine-Tuning on Surgical Phase Recognition (SPR): This table compares the SPR performance of FL-EndoViT and CEN-EndoViT in a few-shot setting across different training set sizes. The reported values represent the accuracy and F1-Score computed per video over the technical repetitions, followed by the mean and standard deviation calcucation across all videos. The bold values indicate weather CEN-EndoViT or FL-EndoViT performs better. Statistically significant differences, determined using the Wilcoxon paired test with $\alpha=0.01$, are marked with an asterisk.
Metric Accuracy F1-Score
Fine-Tuned on TeCNO Stage CEN-EndoViT FL-EndoViT CEN-EndoViT FL-EndoViT
2 Videos S1 $\bm{63.83\%\pm 8.94\%}$* $61.30\%\pm 8.51\%$ $\bm{48.95\%\pm 8.26\%}$* $45.70\%\pm 6.96\%$
[…]
8 Videos S1 $\bm{74.04\%\pm 8.65\%}$* $71.96\%\pm 10.32\%$ $58.76\%\pm 9.90\%$ $\bm{60.62\%\pm 10.02\%}$*
[…]
The results of the SPR fine-tuning experiment are presented in Table 5, while the outcomes of full fine-tuning on the entire dataset are provided in Table 2.
The TeCNO model consists of two stages: Stage S1, which involves feature extraction without temporal components, and Stage S2, where temporal components are incorporated.
In Stage S1, both FL-EndoViT and CEN-EndoViT exhibit similar performance, with only minor differences in F1 scores, as shown in Tables 5 and 2. Model performance varies depending on the videos used for fine-tuning, with either the federated or centralized backbone outperforming the other in certain cases. However, overall, CEN-EndoViT performs better when trained on limited data, whereas with more data available for fine-tuning, FL-EndoViT performs comparably or can even surpass CEN-EndoViT. In Stage S2, which incorporates temporal components, both models perform similarly, with no statistically significant differences, highlighting the robustness of the FL-trained backbone in capturing temporal information.
6 Discussion
In this study, we conducted a detailed comparison between our FL-EndoViT model, incorporating adaptive FedSAM, SWA, and a federated MAE approach, and the established CEN-EndoViT model. We evaluated their performance during the pretraining phase and across multiple downstream surgical tasks, including ATR, SPR, and SSS. Our results demonstrate that FL-trained surgical Foundation Models can match centralized performance in ATR and SPR, while outperforming it in SSS when high-resolution images are used. This highlights their ability to generalize across diverse surgical datasets while preserving privacy and underscores FL’s potential for developing robust and privacy-preserving models in SDS.
6.1 Pretraining Results
Our pretraining experiments highlight the advantages of incorporating adaptive FedSAM into the federated MAE approach. Initially, the basic federated EndoViT model without FedSAM performed suboptimally compared to the centralized model. However, integrating FedSAM led to significant improvements in reconstruction loss and the number of sufficiently reconstructed patches, indicating its effectiveness in handling data heterogeneity. By guiding the optimization process toward flatter minima, FedSAM is theoretically expected to enhance model stability and generalization, as suggested in prior studies.
In addition, extant literature suggests that SWA has the capacity to enhance stabilization in federated learning environments across a range of datasets and institutions. While our results align with these findings, our study does not provide direct numerical evidence for these effects. However, our observations suggest that federated models can be optimized to achieve performance comparable to centralized models through the careful selection of optimization techniques.
6.2 Downstream Tasks Results
Before evaluating the FL-EndoViT backbone across SSS, ATR, and SPR in full-shot and few-shot settings, we conducted an ablation study to determine whether full fine-tuning is necessary or if freezing the backbone while fine-tuning only the task-specific head suffices. The results indicate that a frozen-backbone setting is inadequate for capturing task-specific representations, leading to significantly lower performance. In contrast, models that underwent full fine-tuning exhibited substantial performance gains, highlighting the importance of fine-tuning both the backbone and the head for effective task adaptation. The performance gap observed in federated models further underscores the need for additional adaptation strategies to enhance task-specific generalization. Based on these findings, we fully fine-tuned the backbone in the subsequent evaluations.
We then evaluated FL-EndoViT as a backbone for SSS, ATR, and SPR in both full-shot and few-shot settings, with performance varying based on task type and data resolution.
For SSS, models using the FL-EndoViT backbone demonstrated clear advantages over those with the centralized backbone in high-resolution settings, especially when smaller datasets were used for fine-tuning. The federated approach enabled these models to outperform centralized ones when video data was limited — an essential insight for real-world medical applications, where labeled data is often scarce. This suggests that FL can be particularly beneficial in scenarios with limited annotated data, enabling models to learn effectively across decentralized datasets. However, in low-resolution settings, centralized models consistently outperformed FL-EndoViT models, particularly with fewer videos for training. This underscores the importance of high-quality data for the successful application of FL, as lower-quality data may hinder the performance of FL. The reason for that could be the combination of limited information of local data, amplified heterogeneity, and the sensitivity of Vision Transformers to image quality, which together make it difficult for FL to learn robust global representations from weak or noisy local updates.
For ATR, models using the FL-EndoViT backbone showed superior performance in most cases, particularly when all videos were used for fine-tuning. This highlights the ability of the FL configuration to leverage diverse datasets and improve model performance, as evidenced by the notable improvements in mAP scores over centralized models. These findings align with the idea that FL can enhance generalization by training on more varied data from multiple sources. However, in few-shot settings, the performance gap between federated and centralized models narrowed, suggesting that both configurations perform similarly when data is scarce.
In the context of SPR, models employing FL-EndoViT and CEN-EndoViT demonstrated comparable performance in Stage 1, exhibiting no statistically significant variations across full fine-tuning conditions. This observation indicates that, in the absence of temporal components, both federated and centralized models possess the capacity to extract features with a reasonable degree of effectiveness. However, the observed performance variations in the few-shot setting, contingent on different fine-tuning datasets, underscore the significance of dataset selection and the potential for either model to outperform the other depending on the context.
In Stage 2, which incorporates temporal components, both models showed comparable results, suggesting that the federated model is capable of leveraging temporal information at a similar level to the centralized model. This is notable because it demonstrates that FL-EndoViT can effectively handle temporal dependencies without a significant drop in performance, even when the learned features comes from multiple institutions. These findings indicate that federated backbone models, despite operating in a decentralized setting, can match centralized models in terms of both feature extraction and temporal modeling, further emphasizing the potential of FL in SDS, where data privacy is critical.
7 Conclusion
This study investigates how FL can be leveraged to train Foundation Models that generalize across diverse surgical datasets and tasks while preserving privacy. Our findings demonstrate that federated models, particularly those incorporating adaptive FedSAM in their pretraining, exhibit strong generalization and robustness to data heterogeneity. However, limitations arise, especially when the backbone remains frozen, and fine-tuning both the backbone and the task-specific head is crucial for optimal task-specific performance.
The efficacy of federated foundation models in achieving performance comparable to, or even superior to, centralized models, particularly when fine-tuned on both limited data and large datasets, has been demonstrated. This finding underscores the potential of FL for Foundation Model training, supporting the development of privacy-preserving, robust models for SDS applications.
Additionally, investigating FL in video-based models, rather than solely image-based models, could be a valuable direction for SDS. Surgical procedures are dynamic, requiring models that can effectively capture both spatial and temporal features to improve accuracy and applicability in the operating room. Based on this circumstance, it appears likely that further exploration of FL in video-based models may enhance the capabilities of SDS by integrating collaborative learning with the rich temporal information present in surgical videos. This integration has the potential to drive advancements, leading to the development of more effective and secure AI models for surgical applications.
Disclosure of Interest
JNK declares consulting services for Bioptimus, France; Panakeia, UK; AstraZeneca, UK; and MultiplexDx, Slovakia. Furthermore, he holds shares in StratifAI, Germany, Synagen, Germany, Ignition Lab, Germany; has received an institutional research grant by GSK; and has received honoraria by AstraZeneca, Bayer, Daiichi Sankyo, Eisai, Janssen, Merck, MSD, BMS, Roche, Pfizer, and Fresenius. FRK declares advisory roles for Radical Healthcare, USA; and the Surgical Data Science Collective, USA. MW received speakers fee by KARL STORZ SE & Co KG. SS received speakers fee by Stryker Corporation.
Acknowledgements
This work was co-funded by the European Union through NEARDATA under grant agreement ID 101092644, the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden, and the Federal Ministry of Education and Research of Germany in the programme of “Souverän. Digital. Vernetzt.”, a joint project 6G-life with the project identification number 16KISK001K.
Appendix A Pretraining
Figure 2: Pretraining Results Over 15 Epochs Showing Number of Patches below Different Reconstruction Loss Thresholds: This figure illustrates the number of patches (out of a total of 256) with reconstruction losses falling below thresholds of 0.3, 0.1, 0.05, and 0.01 over 15 epochs, representing reconstruction performance. Each line in the figure tracks the count of patches under specific thresholds, along with SWA (Stochastic Weight Averaging) and observed maximum values. Higher counts of patches below a threshold indicate improved reconstruction accuracy. The figures presented here are derived from the pretraining of the SSS variant. A similar pattern is observed for ATR and SPR.
Appendix B Downstream Tasks
B.1 Few-Shot Videos
In the few-shot experiments, three technical repetitions were performed, with each run trained on distinct, non-overlapping video datasets. The following table lists the video IDs of the subsections used in each run.
Table 6: Video IDs used in the few-shot experiments for SSS, ATR, and SPR across three technical repetitions. Each run utilizes distinct, non-overlapping video datasets, with varying numbers of videos for each task.
SSS 1 Video 2 Videos 4 Videos
Run 1 9 24, 1 18, 48, 25, 1
Run 2 28 25, 48 28, 55, 24, 37
Run 3 43 55, 28 37, 18, 24, 20
ATR 2 Videos 4 Videos 8 Videos
Run 1 27, 5 36, 14, 57, 15 18, 57, 51, 60, 27, 68, 36, 48
Run 2 70, 47 23, 66, 31, 27 2, 36, 5, 47, 48, 8, 6, 18
Run 3 51, 50 42, 31, 6, 52 47, 57, 8, 80, 15, 68, 40, 27
SPR 2 Videos 4 Videos 8 Videos
Run 1 25, 1 8, 28, 13, 4 1, 22, 12, 8, 6, 33, 31, 29
Run 2 33, 11 5, 17, 25, 31 20, 38, 25, 7, 29, 4, 2, 24
Run 3 22, 40 24, 28, 38, 15 18, 28, 17, 2, 40, 36, 7, 34
B.2 Fully Fine-Tuned Experiments
Table 7: Performance Comparison of Surgical Scene Segmentation (Fully Fine-Tuned): The values in the table represents mean Intersection Over Union (IoU) in % over all 1040 images and all classes. The better model of each set is marked as bold. If the one model is significantly better than its counterpart, determined using the Wilcoxon paired test with $\alpha=0.01$, it is marked with an asterisk (*).
Fine-Tuning on 1 Video 2 Videos 4 Videos
Low Res Run 1 FL 26.22% 57.15% 58.02%
CEN 30.86%* 61.16%* 61.57%*
Run 2 FL 48.92% 52.05% 59.51%
CEN 48.92% 54.12%* 61.17%*
Run 3 FL 44.90% 50.18% 57.99%
CEN 47.76%* 53.83%* 59.68%*
High Res Run 1 FL 31.42%* 61.29%* 60.36%*
CEN 13.23% 31.96% 45.63%
Run 2 FL 48.73%* 54.27%* 60.99%*
CEN 32.15% 43.16% 44.84%
Run 3 FL 44.78%* 52.92%* 58.85%*
CEN 30.19% 35.95% 45.55%
Figure 3: Performance Comparison of Surgical Scene Segmentation on High-Resolution Images (Fully Fine-Tuned). The violin plots illustrate the distribution of IoU (Intersection over Union) scores for 1,040 test images. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates a significant improvement in performance with the federated backbone model, orange indicates a significant improvement in performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Figure 4: Performance Comparison of Surgical Scene Segmentation on Low-Resolution Images (Fully Fine-Tuned). The violin plots illustrate the distribution of IoU (Intersection over Union) scores for 1,040 test images. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates a significant improvement in performance with the federated backbone model, orange indicates a significant improvement in performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Table 8: Performance Comparison of Action Triplet Recognition (Fully Fine-Tuned): The values in the table represents mean Average Precision in % over all test videos. The better model of each set is marked as bold. If the one model is significantly better than its counterpart, determined using the Wilcoxon paired test with $\alpha=0.01$, it is marked with an asterisk (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos
Run 1 FL 22.18% 30.22% 34.34%
CEN 21.73% 29.54% 31.79%
Run 2 FL 21.60% 24.53% 32.66%
CEN 23.25% 22.14% 32.98%
Run 3 FL 21.50% 26.98% 30.58%
CEN 22.85% 26.33% 34.70%
Figure 5: Performance Comparison of Action Triplet Recognition (Fully Fine-Tuned). The violin plots show average precision scores distribution for nine test videos. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates significantly better performance with the federated backbone model, orange indicates significantly better performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Figure 6: Performance Comparison of Surgical Phase Recognition (Fully Fine-Tuned, Accuracy). The violin plots illustrate the accuracy score distributions for the two-stage TeCNO approach across 31 videos. The left subfigure depicts the initial stage, Feature Extraction without temporal dimension, while the right subfigure illustrates the subsequent stage with temporal dimension. The left half of the violin plots the federated variant, and the right half plots the centralized variant. The violin plots are further annotated with color-coded labels: purple signifies significantly superior performance with the federated backbone model, orange indicates significantly superior performance with the centralized model, and gray denotes no significant difference, as determined by a Wilcoxon signed-rank test.
Table 9: Performance Comparison of Surgical Phase Recognition (Fully Fine-Tuned, accuracy): The values in the table represents mean accuracy in % over all test videos. The better model of each set is marked as bold. If the one model is significantly better than its counterpart, determined using the Wilcoxon paired test with $\alpha=0.01$, it is marked with an asterisk (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos
S1 - Feature Extraction Run 1 FL 57.01% 64.31% 72.04%
CEN 64.14%* 64.73% 71.51%
Run 2 FL 64.80% 69.02% 72.67%
CEN 63.45% 66.68% 75.60%*
Run 3 FL 62.09% 65.35% 71.17%
CEN 63.90% 68.95%* 75.03%*
S2 - Temporal Component Run 1 FL 71.68% 80.26% 85.87%*
CEN 72.90% 81.32% 82.70%
Run 2 FL 81.01% 76.97% 83.06%
CEN 81.84% 76.38% 85.78%
Run 3 FL 76.35% 82.98% 84.91%
CEN 79.63% 81.55% 84.55%
Figure 7: Performance Comparison of Surgical Phase Recognition (Fully Fine-Tuned, Accuracy). The violin plots illustrate the accuracy score distributions for the two-stage TeCNO approach across 31 videos. The left subfigure depicts the initial stage, Feature Extraction without temporal dimension, while the right subfigure illustrates the subsequent stage with temporal dimension. The left half of the violin plots the federated variant, and the right half plots the centralized variant. The violin plots are further annotated with color-coded labels: purple signifies significantly superior performance with the federated backbone model, orange indicates significantly superior performance with the centralized model, and gray denotes no significant difference, as determined by a Wilcoxon signed-rank test.
Table 10: Performance Comparison of Surgical Phase Recognition (Fully Fine-Tuned, F1-Score): The values in the table represents mean F1 score in % over all test videos. The better model of each set is marked as bold. If the one model is significantly better than its counterpart, determined using the Wilcoxon paired test with $\alpha=0.01$, it is marked with an asterisk (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos
S1 - Feature Extraction Run 1 FL 41.10% 55.50% 60.15%*
CEN 50.82%* 53.42% 52.14%
Run 2 FL 52.54% 54.73%* 61.21%
CEN 51.19% 49.38% 63.43%*
Run 3 FL 43.47% 49.43% 60.49%
CEN 44.83% 54.68%* 60.71%
S2 - Temporal Component Run 1 FL 72.49% 78.48% 82.10%*
CEN 69.34% 78.32% 80.20%
Run 2 FL 78.14% 75.68% 80.27%
CEN 78.65% 76.89% 81.96%
Run 3 FL 75.02% 79.51% 82.27%*
CEN 77.46%* 77.80% 80.02%
B.3 Frozen Backbone Experiments
Figure 8: Performance Comparison of Surgical Scene Segmentation on High-Resolution Images (Frozen Backbone). The violin plots illustrate the distribution of IoU (Intersection over Union) scores for 1,040 test images. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates a significant improvement in performance with the federated backbone model, orange indicates a significant improvement in performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Figure 9: Performance Comparison of Surgical Scene Segmentation on Low-Resolution Images (Frozen Backbone). The violin plots illustrate the distribution of IoU (Intersection over Union) scores for 1,040 test images. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates a significant improvement in performance with the federated backbone model, orange indicates a significant improvement in performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Table 11: Performance Comparison of Surgical Scene Segmentation (Frozen Backbone): The values in the table represents mean Intersection Over Union (IoU) in % over all 1040 images and all classes. The better model of each set is marked as bolt. If the one model is significant better than its counterpart ($\alpha=0.01$) it is marked with a star (*).
Fine-Tuning on 1 Video 2 Videos 4 Videos All Videos
Low Res Run 1 FL 30.47% 56.46% 57.70% 65.13%
CEN 31.87%* 59.69%* 61.13%* 66.80%*
Run 2 FL 47.06% 52.45% 58.43% 64.41%
CEN 48.84%* 53.22%* 61.06%* 66.16%*
Run3 FL 44.80% 52.86% 56.84% 65.32%
CEN 47.27%* 53.42%* 59.05%* 67.68%*
High Res Run 1 FL 32.62% 57.54% 61.04% 67.09%
CEN 35.51%* 59.98%* 62.86%* 68.75%*
Run 2 FL 47.82%* 53.00% 60.25% 66.42%
CEN 47.58% 55.60%* 61.25%* 68.00%*
Run 3 FL 45.33% 48.28% 56.63% 67.36%
CEN 47.81%* 54.63%* 60.06%* 67.76%*
Figure 10: Performance Comparison of Action Triplet Recognition (Frozen Backbone). The violin plots show average precision scores distribution for nine test videos. The left half of the violin plots the federated variant, while the right half plots the centralized variant. The split violins are color-coded: purple indicates significantly better performance with the federated backbone model, orange indicates significantly better performance with the centralized model, and gray indicates no significant difference, as measured by a Wilcoxon signed-rank test.
Table 12: Performance Comparison of Action Triplet Recognition (Frozen Backbone): The values in the table represents mean Average Precision in % over all test videos. The better model of each set is marked as bolt. If the one model is significant better than its counterpart ($\alpha=0.01$) it is marked with a star (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos All Videos
Run 1 FL 17.25% 20.83% 22.40% 26.43%
CEN 19.37% 24.09%* 26.31% 31.28%*
Run 2 FL 16.85% 17.95% 23.63% 27.26%
CEN 19.96%* 20.93%* 27.88% 31.51%*
Run 3 FL 14.97% 20.24% 23.09% 27.00%
CEN 18.30%* 23.89%* 27.12%* 31.21%*
Figure 11: Performance Comparison of Surgical Phase Recognition (Frozen Backbone, Accuracy). The violin plots illustrate the accuracy score distributions for the two-stage TeCNO approach across 31 videos. The left subfigure depicts the initial stage, Feature Extraction without temporal dimension, while the right subfigure illustrates the subsequent stage with temporal dimension. The left half of the violin plots the federated variant, and the right half plots the centralized variant. The violin plots are further annotated with color-coded labels: purple signifies significantly superior performance with the federated backbone model, orange indicates significantly superior performance with the centralized model, and gray denotes no significant difference, as determined by a Wilcoxon signed-rank test.
Figure 12: Performance Comparison of Surgical Phase Recognition (Frozen Backbone, Accuracy). The violin plots illustrate the accuracy score distributions for the two-stage TeCNO approach across 31 videos. The left subfigure depicts the initial stage, Feature Extraction without temporal dimension, while the right subfigure illustrates the subsequent stage with temporal dimension. The left half of the violin plots the federated variant, and the right half plots the centralized variant. The violin plots are further annotated with color-coded labels: purple signifies significantly superior performance with the federated backbone model, orange indicates significantly superior performance with the centralized model, and gray denotes no significant difference, as determined by a Wilcoxon signed-rank test.
Table 13: Performance Comparison of Surgical Phase Recognition (Frozen Backbone): The values in the table represents mean accuracy in % over all test videos. The better model of each set is marked as bolt. If the one model is significant better than its counterpart ($\alpha=0.01$) it is marked with a star (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos All Videos
S1 - Feature Extraction Run1 FL 43.51% 47.00% 49.24% 65.89%
CEN 49.50% 53.65% 60.17% 71.68%
Run2 FL 44.50% 48.26% 50.55% 65.45%
CEN 54.06% 54.43% 61.84% 72.85%
Run 3 FL 47.66% 51.75% 54.92% 65.83%
CEN 56.69% 58.47% 60.89% 72.26%
S2 - Temporal Component Run1 FL 40.17% 39.45% 63.91% 78.06%
CEN 65.52% 64.01% 71.05% 81.63%
Run2 FL 58.53% 52.73% 68.68% 75.60%
CEN 66.05% 73.24% 78.20% 79.38%
Run 3 FL 60.21% 59.20% 65.67% 72.11%
CEN 70.43% 74.17% 74.95% 78.64%
Table 14: Performance Comparison of Surgical Phase Recognition (Frozen Backbone): The values in the table represents mean F1 score in % over all test videos. The better model of each set is marked as bolt. If the one model is significant better than its counterpart ($\alpha=0.01$) it is marked with a star (*).
Fine-Tuning on 2 Videos 4 Videos 8 Videos All Videos
S1 - Feature Extraction Run1 FL 27.84% 34.18% 34.88% 53.75%
CEN 38.11%* 35.83% 41.23%* 58.77%*
Run2 FL 30.63% 29.72% 41.19% 53.29%
CEN 37.31%* 31.96% 46.90%* 60.59%*
Run 3 FL 29.26% 32.66% 40.13% 53.14%
CEN 34.11%* 37.11%* 40.43% 59.77%*
S2 - Temporal Component Run1 FL 56.08% 57.20% 66.65% 74.36%
CEN 68.84%* 67.33%* 69.93% 76.50%
Run2 FL 65.29% 59.11% 69.55% 73.41%
CEN 66.01% 70.28%* 74.81%* 76.84%*
Run 3 FL 64.72% 66.49% 67.84% 69.11%
CEN 68.92%* 70.29%* 71.10%* 72.02%